OpenStudy (anonymous):

at what point on the curve x^2-y^2+x=2 is the tangent line vertical?

OpenStudy (amistre64):

can we take the derivative?

OpenStudy (amistre64):

2x -2yy' +1 =0 -2x -1 ------ = y' y when y=0

OpenStudy (amistre64):

forgot a couple details, but the result is the same

OpenStudy (amistre64):

x^2 + x = 2 x^2 +x -2 =0 x = -4 and x= 2 the point (-4,0) (2,0)

OpenStudy (amistre64):

doh....divide by 2

OpenStudy (amistre64):

(-2,0) and (1,0)

OpenStudy (anonymous):

k, that is what i was going to tell you, thanks a lot

OpenStudy (amistre64):

by all means, when im wrong let me know :)